OpenStudy (anonymous):

integrate (cscx)^3 >_>

OpenStudy (raden):

integration by parts it will work

OpenStudy (raden):

int (cscx)^3 = int (cscx)^2 cscx dx let u=cscx du=.... dv=(cscx)^2 dx v=int(cscx)^2 dx v=..... and next use parts rule : int(udv) = uv-int(vdu)

OpenStudy (raden):

but i was stuck to find int(cscx) dx, @TuringTest ^^ lol

OpenStudy (raden):

i'm not sure int (cscx)dx = ln(sinx)

Directrix (directrix):

@RadEn OpenStudy values the Learning process - not the ‘Give you an answer’ process •Don’t post only answers - guide the asker to a solution. http://openstudy.com/code-of-conduct

OpenStudy (experimentx):

\[ \int \csc ^3xdx = \int (1 + \cot^2 x) \csc x \; dx\]

OpenStudy (anonymous):

is there another trig identity i'm supposed to use? u-substitution isn't working

OpenStudy (anonymous):

There is something of a trick to this... try \[\int \frac{ \sin(x)}{\sin^2(x) } dx\]

OpenStudy (anonymous):

i'm not sure how you got that

OpenStudy (anonymous):

its not (cosx)^2 / (sinx)^3?

OpenStudy (anonymous):

I'm talking about integrating cosecant.

OpenStudy (experimentx):

use the same technique as used here http://en.wikipedia.org/wiki/Integral_of_secant_cubed